OpenStudy (7thgrader):

Which equations have a positive solution? Choose all answers that are correct. A.-w/9 = 8 B. -12 = y/(-3) C.-4 = -2/5 z D.1/7 x -2

OpenStudy (7thgrader):

@mathmath333 @marie123 @henrry123 @Secret-Ninja @shortycme @morganchavez

OpenStudy (7thgrader):

will fan and medal

OpenStudy (anonymous):

idk mann im sorry

OpenStudy (mathmath333):

\(\huge -x=-y\\ \huge x=y\)

OpenStudy (7thgrader):

its ok.

OpenStudy (7thgrader):

????

OpenStudy (7thgrader):

how i learn is you give me the answer but walk me through how to get the answer

OpenStudy (mathmath333):

example if \(-x=1\\ then\\ x=-1\) but if \(-x=-1\\ then\\ x=1\)

OpenStudy (7thgrader):

yhea

OpenStudy (7thgrader):

now what?

OpenStudy (mathmath333):

now see ur options and match them as i figured out two possiblities

OpenStudy (7thgrader):

@mathmath333 can you help me

OpenStudy (mathmath333):

so what hve u figured out

OpenStudy (mathmath333):

ex. 1st option \(\Large\rm \begin{align} \color{black}{\dfrac{-w}{9 }= 8\\~\\ -w= 8\times 9\\~\\ -w= 72\\~\\ w=-72}\end{align}\)

OpenStudy (mathmath333):

so it is negative solution ,so its not positive